Compare the rates of diffusion of (a) O2 and Cl2 (b) CH4 and HCL


1
Expert's answer
2020-11-26T05:07:41-0500

The ratio of rates of diffusion can be found with the molar masses of the molecules:


"\frac{r_{\text{O}_2}}{r_{\text{Cl}_2}}=\sqrt{\frac{\Mu_{\text{Cl}_2}}{\Mu_{\text{O}_2}}}=1.49,\\\space\\

\frac{r_{\text{CH}_4}}{r_{\text{HCl}}}=\sqrt{\frac{\Mu_{\text{HCl}}}{\Mu_{\text{CH}_4}}}=1.51."

As we see, the rate of diffusion of oxygen is greater than that of chlorine; the rate of diffusion of hydrochloric acid is less than the rate of diffusion of methane.